# An analogue of Reed’s conjecture for digraphs

### v2

@@ 4 Hitting every maximum biclique with an acyclic set of vertices
Henceforth, by Claim 4.9, we assume without loss of generality that $\bigcap_{C\in V(\mathcal{H}_{r})}C=\varnothing$. By Lemma 4.8 applied to $S(D)$ with $\mathcal{C}=V(\mathcal{H}_{r})$, we obtain ${\omega}(D)=\omega(S(D))=\frac{2}{3}(\Delta+1)$ and $Y=\bigcup_{C\in\mathcal{C}}C$ admits a partition $Q_{1},\dots,Q_{n}$ into $n\geq 4$ bicliques of $D$ each of cardinality $\frac{1}{2}\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)$ such that $\mathcal{C}\setminus\{Q_{1}\cup Q_{n}\}=\{Q_{i}\cup Q_{i+1}\mid i\in[n-1]\}$.
@@
If $Q_{1}\cup Q_{n}\in\mathcal{C}$, then $D[Y]$ is $\Delta$-diregular, so necessarily $Y=V(D)$ as $D$ is connected, and $D$ is exactly $\overleftrightarrow{C_{n}\circ K_{p}}$. If $n$ is odd, this is a contradiction to $D$ being a counterexample. If $n$ is even, one can arbitrarily pick one representative of each $Q_{i}$ with $i\geq 1$ odd to obtain an acyclic set intersecting every maximum biclique of $D$, a contradiction.
@@
Hence $\mathcal{C}=\{Q_{i}\cup Q_{i+1}\mid i\in[n-1]\}$. Let $D^{\prime}$ be the digraph obtained from $D$ by removing $\bigcup_{i=2}^{n-1}Q_{i}$ and adding every possible arcs between $Q_{1}$ and $Q_{n}$. It is straightforward that $\Delta_{\max}(D)\geq\Delta_{\max}(D^{\prime})$ and that $D^{\prime}$ is smaller than $D$.
@@
Note that ${\omega}(D^{\prime})\geq\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)$ as $Q_{1}\cup Q_{n}$ is a biclique of $D^{\prime}$ of size ${\omega}(D)$. We claim that equality holds, so assume for a contradiction that $D^{\prime}$ contains a biclique $C^{\prime}$ of size ${\omega}(D)+1$. Let $q_{1}=|C^{\prime}\cap Q_{1}|$ and $q_{n}=|C^{\prime}\cap Q_{n}|$. Observe that $q_{1}>0$ and $q_{n}>0$ as $C^{\prime}$ is not a biclique of $D$. Let us fix $v\in C^{\prime}\setminus(Q_{1}\cup Q_{n})$, which exists since $|Q_{1}\cup Q_{n}|<|C^{\prime}|$. Recall that $v$ belongs to a maximum biclique $C$ of $D$, which is necessarily disjoint from $Q_{1}\cup Q_{n}$ (otherwise $C$ would belong to $\mathcal{C}$). If $q_{1}+q_{n}>\frac{1}{3}(\Delta+1)$, then $d^{+}_{D}(v)\geq q_{1}+q_{n}+\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)-1>\Delta$, a contradiction. Hence $q_{1}+q_{n}\leq\frac{1}{3}(\Delta+1)$, which implies $|C^{\prime}\setminus(Q_{1}\cup Q_{n})|=\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)+1-q_{1}-q_{n}\geq\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)+1-\frac{1}{3}(\Delta+1)$. Let $v_{1}$ be any vertex in $C^{\prime}\cap Q_{1}$, then $d^{+}_{D}(v_{1})\geq|C^{\prime}\setminus(Q_{1}\cup Q_{n})|+|Q_{1}\cup Q_{2}|-1\geq 2\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)-\frac{1}{3}(\Delta+1)\geq\Delta+1$, a contradiction.
@@
This shows that ${\omega}(D^{\prime})=\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)$. Note also that $D^{\prime}$ is connected as $D$ is, and vertices in $\bigcup_{i=2}^{n-1}Q_{i}$ do not have neighbours outside $Y$. Since $D$ is a minimum counterexample, either $D^{\prime}=\overleftrightarrow{C_{k}\circ K_{p}}$ or $D^{\prime}$ has an acyclic set $I$ such that ${\omega}(D^{\prime}-I)=\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D^{\prime})-1$. In the former case, we obtain that $D=\overleftrightarrow{C_{k+n-2}\circ K_{p}}$, a contradiction to $D$ being a counter example: either $k+n-2$ is odd or we can easily find an acyclic set intersecting every maximum biclique. Assume the latter case holds. As $I$ intersects every maximum biclique of $D^{\prime}$, in particular it intersects $Q_{1}\cup Q_{n}$. We assume by symmetry that it intersects $Q_{1}$. For every $i\in[n]$, let $v_{i}$ be an arbitrary vertex of $Q_{i}$. If $12\leq i\leq n-$ 1, observe that $N^{+}(v_{i})=N^{-}(v_{i})=(Q_{i-1}\cup Q_{i}\cup Q_{i+1})\setminus\{v_{i}\}$.
@@ 5 Handling vertices with a dense neighbourhood
Theorem 5.1 (King and Reed [24]).
Let $a$ and $\varepsilon$ be constants satisfying $0<\varepsilon<\frac{1}{6}-2\sqrt{a}$. Let $G$ be a graph with maximum degree $\Delta$ satisfying $\omega(G)\leq\frac{2}{3}(\Delta+1)$ and let $v$ be such that $m(v)>(1-a)\binom{n}{2}$. Then
$\chi(G)\leq\max\left(\chi(G-v),(1-\varepsilon)(\Delta+1)\right).$
The goal of this section is to prove Theorem 1.9, which is the direct analogue of Theorem 5.1 for digraphs. We first recall it here for convenience.
See 1.9
@@
Let $G$ be a graph obtained from $K_{n}$ by removing a matching of size $p$. Then $G$ is $(n-p)$-choosable.
@@
Let $D$ be a digraph obtained from $\overleftrightarrow{K_{n}}$ by removing a matching of size $p$. Then $D$ is $(n-p)$-dichoosable.
@@
Let $L$ be any $(n-p)$-list assignment of $D$. By Lemma 5.2, there exists an $L$-colouring $\alpha$ of $S(D)$, because $S(D)$ is obtained from $K_{n}$ by removing a matching of size $p$. As $D$ is obtained from $\overleftrightarrow{S(D)}$ by adding an acyclic set of arcs, $\alpha$ must be an $L$-dicolouring of $D$. ∎
@@
We take $\Delta(a)=\max\left(\frac{1-a}{\sqrt{a}-a},\frac{1-a}{a}\right)$. By directional duality we assume without loss of generality that $m^{+}(v)\geq(1-a)\Delta(\Delta-1)$ and let $N=N^{+}(v)\cup\{v\}$ and $\bar{N}=V(D)\setminus N$. We assume that $|N|=\Delta+1$ for otherwise we add new vertices dominated by $v$. We partition $N$ into three parts as follows:
$N_{1}$ =$=\left\{u\in N\mid|N^{+}(u)\cap\bar{N}|\geq\frac{\Delta}{2}\right\},$
$N_{2}$ =$=\bigg{\{}u\in(N\setminus N_{1})\mid|N^{+}(u)\cap(\bar{N}\cup N_{1})|\geq 2\sqrt{a}\Delta\bigg{\}},\text{and}$
@@
Using the fact that $\Delta\geq\Delta(a)\geq\frac{1-a}{\sqrt{a}-a}$, we thus have $|N_{1}|<2\sqrt{a}\Delta$. By definition, $v\notin N_{1}$ as $N^{+}(v)\subseteq N$. We also have $v\notin N_{2}$ as $N^{+}(v)\cap(\bar{N}\cup N_{1})=N_{1}$, and $|N_{1}|<2\sqrt{a}\Delta$. We will now bound the size of $N_{2}$. Recall that every vertex $u\in N_{1}$ satisfies $2|N^{+}(u)\cap\bar{N}|\geq\Delta\geq d^{-}(u)$. We thus have the following.
@@
which together with $\Delta\geq\Delta(a)\geq\frac{1-a}{a}$ implies $|N_{2}|<2\sqrt{a}\Delta$.
Let $k=\max(\vec{\chi}(D-v),\left\lfloor(1-\varepsilon)(\Delta+1)\right\rfloor)$. Since $v\in N_{3}$, $D-N_{3}$ is $k$-dicolourable, as it is a subdigraph of $D-v$. Let $\alpha$ be such a $k$-dicolouring and $L$ be the list assignment of $D[N_{3}]$ defined as follows
@@
$|L(u)|>k-4\sqrt{a}(\Delta+1)\geq(1-\varepsilon)(\Delta+1)-1-4\sqrt{a}(\Delta+1)\geq\frac{5}{6}(\Delta+1)-1.$
As $|L(u)|$ is an integer, we thus have $|L(u)|\geq\left\lfloor\frac{5}{6}(\Delta+1)\right\rfloor$. Let $M$ be a maximum matching in $\bar{D}[N_{3}]$ (the complementary of $D[N_{3}]$). Let $X$ be the vertices of $N_{3}$ that are not covered $M$. By definition, $X$ is a biclique of $D$ so $|X|\leq\frac{2}{3}(\Delta+1)$. Note that $|N_{3}|=2|M|+|X|$. If $|M|\leq\frac{1}{6}(\Delta+1)$ then $|M|+|X|\leq\frac{5}{6}(\Delta+1)$. If $|M|\geq\frac{1}{6}(\Delta+1)$ then $|M|+|X|=|N_{3}|-|M|\leq\frac{5}{6}(\Delta+1)$.
In both cases, we obtain $|M|+|X|\leq\frac{5}{6}(\Delta+1)$. Since $D[N_{3}]$ is a subdigraph of a digraph obtained from $\overleftrightarrow{K_{|N_{3}|}}$ by removing a matching of size $|M|$, $D[N_{3}]$ is $(|M|+|X|)$-dichoosable by Lemma 5.3. Therefore $D[N_{3}]$ must be $\left\lfloor\frac{5}{6}(\Delta+1)\right\rfloor$-dichoosable, so in particular it is $L$-dicolourable, implying the result. ∎
@@ 6 Proof of the main result
This section is devoted to the proof of our main result, namely Theorem 1.6. We need the following intermediate result proved by the second author [29]. Recall that an oriented graph is a digraph without any digon.
Theorem 6.1 (Picasarri-Arrieta [29]).
Let $D$ be an oriented graph, then $\vec{\chi}(D)\leq\max(2,\Delta_{\min}(D))$.
We are now going to prove Theorem 1.6. Let us first restate it. See 1.6
@@
  • $\gamma_{\varepsilon}=\frac{\varepsilon}{1-\varepsilon}\leq 2\varepsilon$.
We claim that the statement holds for this specific value of $\varepsilon$. Assume this is not the case, and let $D$ be a counterexample with minimum order. For the sake of conciseness let us denote $\tilde{\Delta}(D)$ by $\tilde{\Delta}$, $\Delta_{\max}(D)$ by $\Delta_{\max}$, ${\omega}(D)$ by ${\omega}$, and let $k=\lceil(1-\varepsilon)(\tilde{\Delta}+1)+\varepsilon\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}\rceil$. As $D$ is a minimum counter example, it satisfies $\vec{\chi}(D)>k$ and for every vertex $v\in V$ we have $\vec{\chi}(D-v)\leq k$. This directly implies that every vertex $v$ of $D$ satisfies $\min(d^{-}(v),d^{+}(v))\geq k$, for otherwise a $k$-dicolouring of $D$ can be obtained from any $k$-dicolouring of $D-v$ by choosing for $v$ a colour that is not appearing either in its in-neighbourhood or in its out-neighbourhood. In particular, this implies that $\tilde{\Delta}$ must be close to $\Delta_{\max}$, as we justify in the following claim.
@@
$\Delta_{\max}\leq(1+\gamma_{\varepsilon})\tilde{\Delta}$.
Proof of claim. Assume this is not the case, so there exists a vertex $v$ such that $\max(d^{+}(v),d^{-}(v))>(1+\gamma_{\varepsilon})\tilde{\Delta}$. By directional duality, we assume without loss of generality that $d^{+}(v)\geq d^{-}(v)$. Therefore, we have $\tilde{\Delta}^{2}\geq d^{+}(v)\cdot d^{-}(v)>(1+\gamma_{\varepsilon})\tilde{\Delta}\cdot d^{-}(v)$, which implies $d^{-}(v)<\frac{1}{1+\gamma_{\varepsilon}}\tilde{\Delta}=(1-\varepsilon)\tilde{\Delta}\leq k$, a contradiction. $\lozenge$
With Theorem 4.2 in hands, we can also justify that there exists a significant gap between ${\omega}$ and $\Delta_{\max}$.
@@
${\omega}\leq\frac{2}{3}\left(\Delta_{\max}+1\right)$.
Proof of claim. Assume for a contradiction that ${\omega}>\frac{2}{3}(\Delta_{\max}+1)$. By Theorem 4.2, there exists an acyclic set of vertices $I\subseteq V(D)$ which intersects every maximum biclique of $D$. Among all such sets $I$, we choose one with maximum size. In particular, this implies that every vertex in $V(D)\setminus I$ has at least one in-neighbour and one out-neighbour in $I$. Let $D^{\prime}$ be $D-I$. As $I$ is acyclic and intersects every maximum clique of $D$, we have ${\omega}(D^{\prime})=\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}-1$. The following shows that we also have $\tilde{\Delta}(D^{\prime})\leq\tilde{\Delta}-1$.
@@
$\leq\max\left\{\sqrt{(d^{+}_{D}(v)-1)(d^{-}_{D}(v)-1)}\mid v\in V(D^{\prime})\right\}$
$\leq\max\left\{\sqrt{d^{+}_{D}(v)d^{-}_{D}(v)}-1\mid v\in V(D^{\prime})\right\}$
$\leq\tilde{\Delta}-1,$
where in the first inequality we used the maximality of $I$ and in the second inequality we used that $\sqrt{(p-1)(q-1)}\leq\sqrt{pq}-1$ holds for every pair of real numbers $p,q\geq 1$. We briefly justify it here for completeness.
00 0 $\leq(\sqrt{p}-\sqrt{q})^{2}$
$\Rightarrow$ $-p-q$ $\leq-2\sqrt{pq}$
$\Rightarrow$ $pq-p-q+1$ $\leq pq-2\sqrt{pq}+1$
$\Rightarrow$ $(p-1)(q-1)$ $\leq(\sqrt{pq}-1)^{2}$
$\Rightarrow$ $\sqrt{(p-1)(q-1)}$ $\leq\sqrt{pq}-1$
Since $I$ is acyclic, we have $\vec{\chi}(D)\leq\vec{\chi}(D^{\prime})+1$, as any dicolouring of $D^{\prime}$ can be extended to $D$ by colouring $I$ with an additional colour. Together with the minimality of $D$ we obtain
$\vec{\chi}(D)$ $\leq\left\lceil(1-\varepsilon)(\tilde{\Delta}(D^{\prime})+1)+\varepsilon\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D^{\prime})\right\rceil+1$
$\leq\left\lceil(1-\varepsilon)\tilde{\Delta}+\varepsilon(\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}-1)\right\rceil+1$

### v3

@@ 5 Hitting every maximum biclique with an acyclic set of vertices
Henceforth, by Claim 5.9, we assume without loss of generality that $\bigcap_{C\in V(\mathcal{H}_{r})}C=\varnothing$. By Lemma 5.8 applied to $S(D)$ with $\mathcal{C}=V(\mathcal{H}_{r})$, we obtain $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)=\omega(S(D))=\frac{2}{3}(\Delta+1)$ and $Y=\bigcup_{C\in\mathcal{C}}C$ admits a partition $Q_{1},\dots,Q_{n}$ into $n\geqslant 4$ bicliques of $D$ each of cardinality $\frac{1}{2}\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)$ such that $\mathcal{C}\setminus\{Q_{1}\cup Q_{n}\}=\{Q_{i}\cup Q_{i+1}\mid i\in[n-1]\}$.
@@
If $Q_{1}\cup Q_{n}\in\mathcal{C}$, then $D[Y]$ is $\Delta$-diregular, so necessarily $Y=V(D)$ as $D$ is connected, and $D$ is exactly $\overleftrightarrow{C_{n}\circ K_{p}}$. If $n$ is odd, this is a contradiction to $D$ being a counterexample. If $n$ is even, one can arbitrarily pick one representative of each $Q_{i}$ with $i\geqslant 1$ odd to obtain an acyclic set intersecting every maximum biclique of $D$, a contradiction.
@@
Hence $\mathcal{C}=\{Q_{i}\cup Q_{i+1}\mid i\in[n-1]\}$. Let $D^{\prime}$ be the digraph obtained from $D$ by removing $\bigcup_{i=2}^{n-1}Q_{i}$ and adding every possible arcs between $Q_{1}$ and $Q_{n}$. It is straightforward that $\Delta_{\max}(D)\geqslant\Delta_{\max}(D^{\prime})$ and that $D^{\prime}$ is smaller than $D$.
@@
Note that $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D^{\prime})\geqslant\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)$ as $Q_{1}\cup Q_{n}$ is a biclique of $D^{\prime}$ of size $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)$. We claim that equality holds, so assume for a contradiction that $D^{\prime}$ contains a biclique $C^{\prime}$ of size $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)+1$. Let $q_{1}=|C^{\prime}\cap Q_{1}|$ and $q_{n}=|C^{\prime}\cap Q_{n}|$. Observe that $q_{1}>0$ and $q_{n}>0$ as $C^{\prime}$ is not a biclique of $D$. Let us fix $v\in C^{\prime}\setminus(Q_{1}\cup Q_{n})$, which exists since $|Q_{1}\cup Q_{n}|<|C^{\prime}|$. Recall that $v$ belongs to a maximum biclique $C$ of $D$, which is necessarily disjoint from $Q_{1}\cup Q_{n}$ (otherwise $C$ would belong to $\mathcal{C}$). If $q_{1}+q_{n}>\frac{1}{3}(\Delta+1)$, then $d^{+}_{D}(v)\geqslant q_{1}+q_{n}+\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)-1>\Delta$, a contradiction. Hence $q_{1}+q_{n}\leqslant\frac{1}{3}(\Delta+1)$, which implies $|C^{\prime}\setminus(Q_{1}\cup Q_{n})|=\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)+1-q_{1}-q_{n}\geqslant\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)+1-\frac{1}{3}(\Delta+1)$. Let $v_{1}$ be any vertex in $C^{\prime}\cap Q_{1}$, then $d^{+}_{D}(v_{1})\geqslant|C^{\prime}\setminus(Q_{1}\cup Q_{n})|+|Q_{1}\cup Q_{2}|-1\geqslant 2\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)-\frac{1}{3}(\Delta+1)\geqslant\Delta+1$, a contradiction.
@@
This shows that $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D^{\prime})=\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)$. Note also that $D^{\prime}$ is connected as $D$ is, and vertices in $\bigcup_{i=2}^{n-1}Q_{i}$ do not have neighbours outside $Y$. Since $D$ is a minimum counterexample, either $D^{\prime}=\overleftrightarrow{C_{k}\circ K_{p}}$ or $D^{\prime}$ has an acyclic set $I$ such that $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D^{\prime}-I)=\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D^{\prime})-1$. In the former case, we obtain that $D=\overleftrightarrow{C_{k+n-2}\circ K_{p}}$, a contradiction to $D$ being a counterexample: either k+n2k+n-2 is odd or we can easily find an acyclic set intersecting every maximum biclique. Assume the latter case holds. As $I$ intersects every maximum biclique of $D^{\prime}$, in particular it intersects $Q_{1}\cup Q_{n}$. We assume by symmetry that it intersects $Q_{1}$. For every $i\in[n]$, let $v_{i}$ be an arbitrary vertex of $Q_{i}$. If $2\leqslant i\leqslant n-1$, observe that $N^{+}(v_{i})=N^{-}(v_{i})=(Q_{i-1}\cup Q_{i}\cup Q_{i+1})\setminus\{v_{i}\}$.
@@ 6 Handling vertices with a dense neighbourhood
Theorem 6.1 (King and Reed [26, Theorem 6]).
Let $a$ and $\varepsilon$ be constants satisfying $0<\varepsilon<\frac{1}{6}-2\sqrt{a}$. Let $G$ be a graph with maximum degree $\Delta$ satisfying $\omega(G)\leqslant\frac{2}{3}(\Delta+1)$ and let $v$ be such that $m(v)>(1-a)\binom{\Delta}{2}$. Then
$\chi(G)\leqslant\max\left(\chi(G-v),(1-\varepsilon)(\Delta+1)\right).$
The goal of this section is to prove Theorem 1.10, which is a direct analogue of Theorem 6.1 for digraphs. We first recall it here for convenience.
See 1.10
@@
Let $G$ be a graph obtained from $K_{n}$ by removing a matching of size $p$. Then $G$ is (np)(n-p)-choosable.
@@
Let $D$ be a digraph obtained from $\overleftrightarrow{K_{n}}$ by removing a matching of size $p$. Then $D$ is (np)(n-p)-dichoosable.
@@
Let $L$ be any (np)(n-p)-list assignment of $D$. By Lemma 6.2, there exists an $L$-colouring $\alpha$ of $S(D)$, because $S(D)$ is obtained from $K_{n}$ by removing a matching of size $p$. As $D$ is obtained from $\overleftrightarrow{S(D)}$ by adding an acyclic set of arcs, $\alpha$ must be an $L$-dicolouring of $D$. ∎
@@
We take $\Delta(a)=\max\left(\frac{1-a}{\sqrt{a}-a},\frac{1-a}{a}\right)$. By directional duality we assume without loss of generality that $m^{+}(v)\geqslant(1-a)\Delta(\Delta-1)$ and let $N=N^{+}(v)\cup\{v\}$ and $\bar{N}=V(D)\setminus N$. We assume that $|N|=\Delta+1$ for otherwise we add new vertices dominated by $v$. We partition $N$ into three parts as follows:
$N_{1}$ $=\left\{u\in N\mid|N^{+}(u)\cap\bar{N}|\geqslant\frac{\Delta}{2}\right\},$
$N_{2}$ $=\bigg{\{}u\in(N\setminus N_{1})\mid|N^{+}(u)\cap(\bar{N}\cup N_{1})|\geqslant 2\sqrt{a}\Delta\bigg{\}},\text{and}$
@@
Using the fact that $\Delta\geqslant\Delta(a)\geqslant\frac{1-a}{\sqrt{a}-a}$, we thus have $|N_{1}|<2\sqrt{a}\Delta$. By definition, $v\notin N_{1}$ as $N^{+}(v)\subseteq N$. We also have $v\notin N_{2}$ as $N^{+}(v)\cap(\bar{N}\cup N_{1})=N_{1}$, and $|N_{1}|<2\sqrt{a}\Delta$. We will now bound the size of $N_{2}$. Recall that every vertex $u\in N_{1}$ satisfies $2|N^{+}(u)\cap\bar{N}|\geqslant\Delta\geqslant d^{-}(u)$. We thus have the following.
@@
which together with $\Delta\geqslant\Delta(a)\geqslant\frac{1-a}{a}$ implies $|N_{2}|<2\sqrt{a}\Delta$.
Let $k=\max(\vec{\chi}(D-v),\left\lfloor(1-\varepsilon)(\Delta+1)\right\rfloor)$. Since $v\in N_{3}$, $D-N_{3}$ is $k$-dicolourable, as it is a subdigraph of DvD-v. Let $\alpha$ be such a $k$-dicolouring and $L$ be the list assignment of $D[N_{3}]$ defined as follows
@@
$|L(u)|>k-4\sqrt{a}(\Delta+1)\geqslant(1-\varepsilon)(\Delta+1)-1-4\sqrt{a}(\Delta+1)\geqslant\frac{5}{6}(\Delta+1)-1.$
As $|L(u)|$ is an integer, we thus have $|L(u)|\geqslant\left\lfloor\frac{5}{6}(\Delta+1)\right\rfloor$. Let $M$ be a maximum matching in $\bar{D}[N_{3}]$ (the complement of $D[N_{3}]$). Let $X$ be the vertices of $N_{3}$ that are not covered $M$. By definition, $X$ is a biclique of $D$ so $|X|\leqslant\frac{2}{3}(\Delta+1)$. Note that $|N_{3}|=2|M|+|X|$. If $|M|\leqslant\frac{1}{6}(\Delta+1)$ then $|M|+|X|\leqslant\frac{5}{6}(\Delta+1)$. If $|M|\geqslant\frac{1}{6}(\Delta+1)$ then $|M|+|X|=|N_{3}|-|M|\leqslant\frac{5}{6}(\Delta+1)$.
In both cases, we obtain $|M|+|X|\leqslant\frac{5}{6}(\Delta+1)$. Since $D[N_{3}]$ is a subdigraph of a digraph obtained from $\overleftrightarrow{K_{|N_{3}|}}$ by removing a matching of size $|M|$, $D[N_{3}]$ is $(|M|+|X|)$-dichoosable by Lemma 6.3. Therefore $D[N_{3}]$ must be $\left\lfloor\frac{5}{6}(\Delta+1)\right\rfloor$-dichoosable, so in particular it is $L$-dicolourable, implying the result. ∎
@@ 7 Proof of the main result
This section is devoted to the proof of our main result, namely Theorem 1.7. We again make use of Theorem 3.2, that we recall here for convenience. Recall that an oriented graph is a digraph without any digon.
Theorem 7.1 (Picasarri-Arrieta [32, Corollary 8]).
Every oriented graph $D$ satisfies
$\vec{\chi}(D)\leqslant\max(2,\Delta_{\min}(D)).$
We are now going to prove Theorem 1.7. Let us first restate it. See 1.7
@@
  • $\gamma_{\varepsilon}=\frac{\varepsilon}{1-\varepsilon}\leqslant 2\varepsilon$.
We claim that the statement holds for this specific value of $\varepsilon$. Assume this is not the case, and let $D$ be a counterexample with minimum order. For the sake of conciseness let us denote $\tilde{\Delta}(D)$ by $\tilde{\Delta}$, $\Delta_{\max}(D)$ by $\Delta_{\max}$, $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)$ by $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}$, and let $k=\lceil(1-\varepsilon)(\tilde{\Delta}+1)+\varepsilon\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}\rceil$. As $D$ is a minimum counterexample, it satisfies $\vec{\chi}(D)>k$ and for every vertex $v\in V$ we have $\vec{\chi}(D-v)\leqslant k$. This directly implies that every vertex $v$ of $D$ satisfies $\min(d^{-}(v),d^{+}(v))\geqslant k$, for otherwise a $k$-dicolouring of $D$ can be obtained from any $k$-dicolouring of DvD-v by choosing for $v$ a colour that is not appearing either in its in-neighbourhood or in its out-neighbourhood. In particular, this implies that $\tilde{\Delta}$ must be close to $\Delta_{\max}$, as we justify in the following claim.
@@
$\Delta_{\max}\leqslant(1+\gamma_{\varepsilon})\tilde{\Delta}$.
Proof of claim. Assume this is not the case, so there exists a vertex $v$ such that $\max(d^{+}(v),d^{-}(v))>(1+\gamma_{\varepsilon})\tilde{\Delta}$. By directional duality, we assume without loss of generality that $d^{+}(v)\geqslant d^{-}(v)$. Therefore, we have $\tilde{\Delta}^{2}\geqslant d^{+}(v)\cdot d^{-}(v)>(1+\gamma_{\varepsilon})\tilde{\Delta}\cdot d^{-}(v)$, which implies $d^{-}(v)<\frac{1}{1+\gamma_{\varepsilon}}\tilde{\Delta}=(1-\varepsilon)\tilde{\Delta}\leqslant k$, a contradiction. $\lozenge$
With Theorem 5.2 in hands, we can also justify that there exists a significant gap between $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}$ and $\Delta_{\max}$.
@@
$\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}\leqslant\frac{2}{3}\left(\Delta_{\max}+1\right)$.
Proof of claim. Assume for a contradiction that $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}>\frac{2}{3}(\Delta_{\max}+1)$. By Theorem 5.2, there exists an acyclic set of vertices $I\subseteq V(D)$ which intersects every maximum biclique of $D$. Among all such sets $I$, we choose one with maximum size. In particular, this implies that every vertex in $V(D)\setminus I$ has at least one in-neighbour and one out-neighbour in $I$. Let $D^{\prime}$ be DID-I. As $I$ is acyclic and intersects every maximum clique of $D$, we have $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D^{\prime})=\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}-1$. The following shows that we also have $\tilde{\Delta}(D^{\prime})\leqslant\tilde{\Delta}-1$.
@@
$\leqslant\max\left\{\sqrt{(d^{+}_{D}(v)-1)(d^{-}_{D}(v)-1)}\mid v\in V(D^{\prime})\right\}$
$\leqslant\max\left\{\sqrt{d^{+}_{D}(v)d^{-}_{D}(v)}-1\mid v\in V(D^{\prime})\right\}$
$\leqslant\tilde{\Delta}-1,$
where in the first inequality we used the maximality of $I$ and in the second inequality we used that $\sqrt{(p-1)(q-1)}\leqslant\sqrt{pq}-1$ holds for every pair of real numbers $p,q\geqslant 1$. We briefly justify it here for completeness.
$0$ $\leqslant(\sqrt{p}-\sqrt{q})^{2}$
$\Rightarrow$ $-p-q$ $\leqslant-2\sqrt{pq}$
$\Rightarrow$ $pq-p-q+1$ $\leqslant pq-2\sqrt{pq}+1$
$\Rightarrow$ $(p-1)(q-1)$ $\leqslant(\sqrt{pq}-1)^{2}$
$\Rightarrow$ $\sqrt{(p-1)(q-1)}$ $\leqslant\sqrt{pq}-1$
Since $I$ is acyclic, we have $\vec{\chi}(D)\leqslant\vec{\chi}(D^{\prime})+1$, as any dicolouring of $D^{\prime}$ can be extended to $D$ by colouring $I$ with an additional colour. Together with the minimality of $D$ we obtain
$\vec{\chi}(D)$ $\leqslant\left\lceil(1-\varepsilon)(\tilde{\Delta}(D^{\prime})+1)+\varepsilon\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D^{\prime})\right\rceil+1$
$\leqslant\left\lceil(1-\varepsilon)\tilde{\Delta}+\varepsilon(\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}-1)\right\rceil+1$